OpenStudy (anonymous):

6+5|4-3to the second|=? A. –19 B. 11 C. 31 D. 55

OpenStudy (muzzack):

Rearrange : Rearrange the equation by subtracting what is to the right of the equal sign from both sides of the equation : 6-9*w-(24)=0 Step by step solution : Step 1 : Pulling out like terms : 1.1 Pull out like factors : -18 - 9w = -9 • (w + 2) Equation at the end of step 1 : -9 • (w + 2) = 0 Step 2 : Solve -9•(w+2) = 0 Equations which are never true : 2.1 Solve : -9 = 0 This equation has no solution. A a non-zero constant never equals zero. Solving a Single Variable Equation : 2.2 Solve : w+2 = 0 Subtract 2 from both sides of the equation : w = -2 One solution was found : w = -2